OpenStudy (anonymous):

use the following information for this problem: p=d(q)=240-6q p=s(q)=q^2 = 28q find: (q,p) market equilibrium point

OpenStudy (anonymous):

hmm this is a hard one i dont know what market equilibrium point mean in math

OpenStudy (anonymous):

neither do I haha

OpenStudy (anonymous):

lol ok ill google it

OpenStudy (anonymous):

the answer choices are (9, 105) or (5, 210) or (6,195) or (6, 204)

OpenStudy (anonymous):

the above equation is supposed to be p=s(q)=q^2+28q not =28q my apologies

OpenStudy (anonymous):

ohhh that is better

OpenStudy (anonymous):

oohhh i think it is simple like the 2 equations have to equal each other

OpenStudy (anonymous):

what answer did you get

OpenStudy (anonymous):

you have to let the equation equal to each other and find your q value :)

OpenStudy (anonymous):

240-6q=q^2+28 q^2+6q-212=0

OpenStudy (anonymous):

ohh q is ugly numbers wait a sec

OpenStudy (anonymous):

@hamza_b23 what answer did you get

OpenStudy (anonymous):

ohh so q is 6 and -40 but that isnt the answer yet give a few more secs to finish off

OpenStudy (anonymous):

okay cool thanks

OpenStudy (anonymous):

i wld say its 6 and 204

OpenStudy (anonymous):

ohh yes it is lol

OpenStudy (anonymous):

really? thankyou so much

OpenStudy (anonymous):

@samjordon do you know how to fin the producers surplus?

OpenStudy (anonymous):

lol i can google it lol

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20081030102159AAEgh4Y i found smth on it that i guess i can follow

OpenStudy (anonymous):

yea i can kind of follow it but were you able to calculate it? @samjordon

OpenStudy (anonymous):

ohh of these formulas ok give me a sec

OpenStudy (anonymous):

I laso need to find the consumers surplus

OpenStudy (anonymous):

ohhh sheg u knwo this stuff no?

OpenStudy (anonymous):

Market equibrium would be when supply is equal to demand

OpenStudy (anonymous):

lol yeah i know lil bit not too much @samjordon

OpenStudy (anonymous):

hahahah ok ill try this out but do u have choices for answers?

OpenStudy (anonymous):

yea i do for consumers surplus my answer choices are 410, 108, 175, 201

OpenStudy (anonymous):

hey in supply equation are u having q^2 = 28q????

OpenStudy (anonymous):

no i typed it wong it is q^2+28q @sheg

OpenStudy (anonymous):

ok now equate these two equations and solve for q

OpenStudy (anonymous):

to get consumers surplus?

OpenStudy (anonymous):

when demand is equal to supply what it will be called as @taffytwink

OpenStudy (anonymous):

what answer do you get @sheg

OpenStudy (anonymous):

i have not solved it you solve it and tell me

OpenStudy (anonymous):

i cannot figure it out lol

OpenStudy (anonymous):

it is \[q^2 +28q = 240 - 6q\] \[q^2 +28q + 6q - 240=0\] \[q^2 + 34q -240=0\] now solve it

OpenStudy (anonymous):

how do I solve it?

OpenStudy (anonymous):

give him the answer sheg :)

OpenStudy (anonymous):

i wont @samjordon

OpenStudy (anonymous):

lol hahahaha

OpenStudy (anonymous):

i am heading to bed nite guys

OpenStudy (anonymous):

@samjordon you are going to bed?! thanks for all the help

OpenStudy (anonymous):

@sheg i still cant figure it out

OpenStudy (anonymous):

hey buddy you want to determine the market equilibrium price and quantity right